# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v2

@@ Appendix E Full results
space_ga 7% -6% 12% 0.86 2.17 0.23 -17% -18% 0.30
space_ga 3% -5% 8% 0.87 2.14 0.23 -14% -15% 0.41
space_ga 2% -5% 5% 0.88 2.23 0.24 -9% -10% 0.41
space_ga 7% -5% 9% 0.86 2.21 0.24 -7% -7% 0.40
space_ga -1% -3% 0% 0.91 2.39 0.26 -14% -18% 0.47
cpu_activity -8% -5% -8% 0.93 0.90 0.17 -36% -46% 0.93
cpu_activity -8% -6% -4% 0.92 0.89 0.17 2% -23% 0.87
cpu_activity -30% -3% -43% 0.94 1.33 0.25 -36% -80% 0.73
cpu_activity -27% -3% -41% 0.94 1.21 0.23 -13% -49% 0.74
cpu_activity -6% -6% -3% 0.91 1.00 0.19 -65% -70% 0.67
naval_propulsion_plant 55% -22% 48% 0.77 1.29 0.38 -18% -50% 0.64
naval_propulsion_plant 55% -10% 42% 0.89 1.52 0.45 -66% -158% 0.79
naval_propulsion_plant 54% -15% 44% 0.84 1.44 0.42 -29% -73% 0.74
naval_propulsion_plant 67% -16% 50% 0.83 1.29 0.38 -56% -142% 0.74
naval_propulsion_plant 57% -20% 48% 0.78 1.38 0.41 -14% -42% 0.66
miami_housing 20% -6% 22% 0.92 1.33 0.17 -13% -26% 0.78
miami_housing 32% -16% 31% 0.79 0.87 0.11 -41% -50% 0.48
miami_housing -4% -5% -1% 0.92 1.39 0.17 -38% -51% 0.67
miami_housing 37% -14% 36% 0.84 0.88 0.11 -21% -34% 0.79
miami_housing 16% -7% 16% 0.89 1.30 0.16 -20% -26% 0.65
kin8nm 5% -9% 13% 0.86 2.48 0.49 -10% -12% 0.32
kin8nm 4% -8% 9% 0.87 2.58 0.51 -10% -12% 0.33
kin8nm 5% -8% 12% 0.87 2.62 0.51 -13% -15% 0.30
kin8nm 8% -11% 17% 0.84 2.41 0.47 -8% -8% 0.30
kin8nm 1% -5% 6% 0.90 2.66 0.52 -14% -18% 0.37
concrete_compressive_strength 4% -11% 15% 0.79 1.88 0.38 -24% -36% 0.46
concrete_compressive_strength -3% -17% 12% 0.78 1.74 0.35 -26% -36% 0.42
concrete_compressive_strength 4% -12% 13% 0.81 1.83 0.37 -30% -47% 0.56
concrete_compressive_strength 7% -11% 15% 0.83 1.81 0.37 -19% -31% 0.58
concrete_compressive_strength -7% -18% 9% 0.80 1.87 0.38 -76% -99% 0.45
cars -18% 1% -39% 0.89 1.29 0.22 -16% -15% 0.81
cars -18% -11% -11% 0.84 1.39 0.23 -23% -30% 0.72
cars -33% -6% -54% 0.84 1.63 0.28 -59% -59% 0.60
cars -28% -12% -41% 0.80 1.27 0.21 -18% -10% 0.59
cars -19% -14% -29% 0.79 1.48 0.25 -49% -42% 0.59
energy_efficiency -11% 13% -35% 0.91 0.50 0.14 -7% -13% 0.98
energy_efficiency 0% -7% 0% 0.83 0.49 0.14 -8% -20% 0.95
energy_efficiency -9% -14% 0% 0.76 0.47 0.14 -112% -137% 0.93
energy_efficiency -11% 4% -30% 0.87 0.59 0.17 -38% -76% 0.96
energy_efficiency -46% 2% -114% 0.89 0.52 0.15 -198% -176% 0.86
california_housing 30% -7% 28% 0.92 2.09 0.50 -11% -13% 0.64
california_housing 28% -9% 28% 0.87 2.02 0.48 -14% -18% 0.50
california_housing 13% -6% 17% 0.92 2.24 0.53 -19% -23% 0.64
california_housing 19% -5% 20% 0.93 2.23 0.53 -12% -16% 0.64
california_housing 16% -6% 18% 0.93 2.35 0.56 -11% -16% 0.60
airfoil_self_noise -7% -18% 10% 0.77 2.02 0.43 -27% -51% 0.36
airfoil_self_noise -26% -18% -11% 0.80 2.21 0.47 -171% -229% 0.30
airfoil_self_noise 25% -20% 36% 0.74 1.28 0.27 -28% -38% 0.41
airfoil_self_noise -22% -24% -2% 0.73 2.26 0.48 -54% -86% 0.13
airfoil_self_noise -7% -18% 10% 0.79 2.02 0.43 -44% -58% 0.43
QSAR_fish_toxicity 11% -8% 15% 0.85 2.37 0.37 -11% -10% 0.40
QSAR_fish_toxicity 13% -8% 18% 0.82 2.02 0.31 -1% -5% 0.47
QSAR_fish_toxicity 19% -7% 22% 0.84 1.90 0.29 -3% -3% 0.48
QSAR_fish_toxicity -3% -15% 12% 0.76 2.02 0.32 -24% -27% 0.25
QSAR_fish_toxicity 8% -9% 14% 0.83 2.29 0.36 -3% -5% 0.36
@@
space_ga -12% -16% 4% 0.68 1.50 0.16 -41% -40% 0.15
space_ga -21% -17% -11% 0.64 1.41 0.15 -85% -61% -0.21
space_ga -5% -14% 6% 0.65 1.30 0.14 -21% -20% 0.37
space_ga -9% -13% 4% 0.70 1.50 0.16 -28% -25% 0.35
space_ga -9% -13% 1% 0.73 1.49 0.16 -25% -26% 0.38
cpu_activity -38% 14% -89% 0.87 0.75 0.14 14% 1% 0.92
cpu_activity -60% 32% -240% 0.94 1.08 0.20 -20% -17% 0.78
cpu_activity -67% 36% -355% 0.95 1.50 0.28 -30% -34% 0.72
cpu_activity -24% -2% -28% 0.68 0.39 0.07 -3% -16% 0.78
cpu_activity -63% 28% -251% 0.85 1.13 0.21 -37% -49% 0.54
naval_propulsion_plant -41% 0% -70% 0.63 0.85 0.25 -62% -68% -0.21
naval_propulsion_plant -53% -13% -87% 0.82 1.60 0.47 -141% -163% 0.50
naval_propulsion_plant -48% -3% -85% 0.72 1.23 0.36 -99% -94% 0.16
naval_propulsion_plant -88% -61% -210% 0.29 1.10 0.33 -215% -562% -0.85
naval_propulsion_plant -33% 4% -55% 0.70 1.04 0.31 -24% -27% 0.27
miami_housing -48% 22% -148% 0.93 1.33 0.17 -21% -20% 0.74
miami_housing -54% 27% -202% 0.89 1.33 0.17 -41% -48% 0.53
miami_housing -19% 11% -40% 0.84 0.75 0.09 -18% -19% 0.75
miami_housing -17% 6% -32% 0.82 0.69 0.09 -23% -22% 0.77
miami_housing -28% 1% -39% 0.76 0.82 0.10 -24% -32% 0.57
kin8nm -7% -5% -3% 0.75 1.67 0.33 -12% -10% 0.39
kin8nm -11% -6% -8% 0.77 1.80 0.35 -20% -22% 0.36
kin8nm -19% -4% -19% 0.77 1.88 0.37 -20% -19% 0.31
kin8nm -4% -2% -5% 0.77 1.68 0.33 -3% 0% 0.43
kin8nm -15% -9% -8% 0.78 1.79 0.35 -26% -24% 0.36
concrete_compressive_strength -15% 5% -26% 0.74 1.53 0.31 -9% -6% 0.51
concrete_compressive_strength -5% -38% 34% 0.40 0.57 0.12 -10% -10% 0.46
concrete_compressive_strength -32% 21% -80% 0.84 1.69 0.34 -5% -9% 0.60
concrete_compressive_strength 2% -53% 54% 0.27 0.39 0.08 -20% -23% 0.52
concrete_compressive_strength -25% 0% -33% 0.76 1.37 0.28 -15% -13% 0.59
cars -50% 38% -176% 0.93 1.62 0.27 -36% -38% 0.75
cars -44% 49% -169% 0.93 1.55 0.26 -3% -5% 0.75
cars -53% 32% -232% 0.94 2.04 0.34 -34% -33% 0.70
cars -49% 44% -223% 0.89 1.59 0.27 -32% -15% 0.48
cars -57% 51% -257% 0.93 1.90 0.32 -32% -37% 0.63
energy_efficiency -35% 4% -68% 0.81 0.43 0.13 -8% 1% 0.97
energy_efficiency -12% 25% -47% 0.60 0.27 0.08 8% 2% 0.92
energy_efficiency -19% -14% -5% 0.57 0.31 0.09 -1% -32% 0.93
energy_efficiency -18% 7% -29% 0.83 0.50 0.14 -8% 6% 0.96
energy_efficiency 20% 26% -3% 0.51 0.12 0.03 14% 18% 0.98
california_housing -16% 1% -20% 0.88 1.76 0.42 -7% -3% 0.63
california_housing -14% 4% -25% 0.83 1.66 0.40 -4% -3% 0.54
california_housing -9% 0% -10% 0.89 1.79 0.43 -1% 0% 0.68
california_housing -20% 5% -30% 0.90 1.91 0.45 -2% -1% 0.65
california_housing -27% 10% -52% 0.89 2.09 0.50 -8% -9% 0.55
airfoil_self_noise 3% -1% 0% 0.76 1.22 0.26 -3% 1% 0.54
airfoil_self_noise 27% -7% 26% 0.77 0.84 0.18 -20% -11% 0.84
airfoil_self_noise 8% -7% 14% 0.62 0.63 0.14 -22% -13% 0.31
airfoil_self_noise -9% 1% -12% 0.76 1.44 0.31 -23% -20% 0.44
airfoil_self_noise -20% 6% -33% 0.83 1.62 0.35 -9% -1% 0.62
QSAR_fish_toxicity -17% -1% -22% 0.75 1.89 0.29 -21% -28% 0.28
QSAR_fish_toxicity -24% 9% -45% 0.79 1.97 0.31 -8% -9% 0.44
QSAR_fish_toxicity -19% 0% -23% 0.70 1.56 0.24 -6% -8% 0.39
QSAR_fish_toxicity -17% 16% -41% 0.76 1.72 0.27 -21% -17% 0.28
QSAR_fish_toxicity -9% 8% -18% 0.81 1.98 0.31 -10% -9% 0.32
Appendix F Example of learning with non-normality

### v3

@@ Appendix E Full conformalized results
space_ga -7% -2% -6% 0.94 3.08 0.33 -18% -19% 0.29
space_ga -12% 3% -18% 0.97 3.41 0.37 -7% -7% 0.40
space_ga -13% 0% -16% 0.95 3.05 0.33 -13% -11% 0.42
space_ga -13% 0% -18% 0.96 3.21 0.35 -8% -8% 0.41
space_ga -10% 0% -11% 0.96 2.91 0.32 -15% -18% 0.46
cpu_activity -18% -3% -23% 0.95 1.03 0.19 -40% -56% 0.92
cpu_activity -12% -6% -11% 0.91 1.04 0.20 -15% -44% 0.81
cpu_activity -5% -5% -2% 0.92 1.00 0.19 -43% -59% 0.75
cpu_activity -13% -2% -14% 0.96 0.97 0.18 2% -19% 0.87
cpu_activity -30% -1% -42% 0.96 1.31 0.25 -10% -43% 0.76
naval_propulsion_plant -19% 0% -23% 0.98 3.17 0.94 -63% -158% 0.71
naval_propulsion_plant 11% -5% 13% 0.94 2.24 0.66 -31% -78% 0.73
naval_propulsion_plant 14% -2% 14% 0.97 2.26 0.67 -61% -147% 0.80
naval_propulsion_plant 24% -9% 25% 0.89 1.99 0.59 -13% -42% 0.66
naval_propulsion_plant -10% -2% -11% 0.95 2.75 0.81 -19% -52% 0.64
miami_housing -15% -2% -17% 0.94 1.86 0.23 -11% -25% 0.79
miami_housing -44% 5% -87% 0.96 2.39 0.30 -43% -64% 0.47
miami_housing -41% 0% -69% 0.96 2.30 0.29 -27% -42% 0.72
miami_housing -35% -2% -52% 0.96 2.18 0.27 -21% -33% 0.79
miami_housing -31% 1% -48% 0.95 2.24 0.28 -22% -30% 0.64
kin8nm -15% -1% -19% 0.96 3.55 0.69 -8% -8% 0.29
kin8nm -14% 2% -19% 0.96 3.43 0.67 -10% -14% 0.32
kin8nm -11% 0% -13% 0.94 3.20 0.63 -9% -12% 0.32
kin8nm -14% 0% -18% 0.95 3.30 0.65 -14% -18% 0.37
kin8nm -14% 2% -19% 0.96 3.42 0.67 -13% -16% 0.30
concrete_compressive_strength -20% 4% -35% 0.96 3.01 0.61 -30% -43% 0.41
concrete_compressive_strength -16% -8% -10% 0.92 3.08 0.62 -72% -94% 0.47
concrete_compressive_strength -18% -3% -24% 0.96 2.99 0.60 -21% -33% 0.56
concrete_compressive_strength -25% -4% -34% 0.94 2.86 0.58 -26% -36% 0.41
concrete_compressive_strength -22% 2% -31% 0.96 3.05 0.62 -35% -50% 0.53
cars -24% -6% -27% 0.92 1.99 0.34 -22% -23% 0.73
cars -49% 4% -103% 0.97 2.70 0.46 -20% -17% 0.80
cars -14% -5% -12% 0.93 2.02 0.34 -59% -56% 0.60
cars -36% -6% -54% 0.93 2.44 0.41 -37% -37% 0.66
cars -37% -6% -68% 0.89 2.12 0.36 -50% -25% 0.34
energy_efficiency 2% -3% 5% 0.97 0.62 0.18 -18% -32% 0.98
energy_efficiency -42% -6% -62% 0.94 0.85 0.25 -159% -104% 0.90
energy_efficiency -4% -6% 2% 0.93 0.69 0.20 -8% -20% 0.95
energy_efficiency -29% -4% -41% 0.94 0.79 0.23 -39% -82% 0.95
energy_efficiency -75% 0% -316% 1.00 2.61 0.76 -122% -136% 0.92
california_housing -2% -2% 0% 0.96 2.88 0.69 -11% -16% 0.60
california_housing -9% -2% -8% 0.96 2.90 0.69 -19% -23% 0.64
california_housing -5% -2% -4% 0.97 3.02 0.72 -11% -13% 0.64
california_housing -1% -3% 1% 0.93 2.77 0.66 -15% -18% 0.50
california_housing -10% -1% -10% 0.97 3.07 0.73 -13% -17% 0.64
airfoil_self_noise -33% -3% -46% 0.97 3.62 0.77 -163% -218% 0.34
airfoil_self_noise -36% -2% -56% 0.93 3.42 0.73 -27% -36% 0.42
airfoil_self_noise -31% -5% -40% 0.92 3.42 0.73 -50% -83% 0.17
airfoil_self_noise -32% -2% -45% 0.97 3.91 0.83 -45% -57% 0.43
airfoil_self_noise -22% -2% -30% 0.94 3.29 0.70 -27% -51% 0.36
QSAR_fish_toxicity 2% 1% 0% 0.98 4.12 0.64 4% 2% 0.51
QSAR_fish_toxicity -10% -4% -7% 0.93 3.43 0.53 -12% -12% 0.40
QSAR_fish_toxicity -7% 1% -11% 0.93 2.87 0.45 -6% -8% 0.45
QSAR_fish_toxicity -17% -7% -16% 0.86 2.99 0.47 -23% -26% 0.25
QSAR_fish_toxicity 0% -3% -1% 0.90 2.63 0.41 -4% -6% 0.34
@@
space_ga -25% -3% -32% 0.95 3.45 0.37 -41% -40% 0.15
space_ga -22% -3% -27% 0.96 3.58 0.39 -28% -25% 0.35
space_ga -45% -2% -82% 0.95 4.52 0.49 -85% -61% -0.21
space_ga -13% -2% -13% 0.95 3.12 0.34 -21% -20% 0.37
space_ga -20% -6% -18% 0.92 3.07 0.33 -25% -26% 0.38
cpu_activity -22% -2% -32% 0.95 1.14 0.21 14% 1% 0.92
cpu_activity -44% 2% -80% 0.95 1.40 0.27 -30% -34% 0.72
cpu_activity -49% 1% -100% 0.94 1.44 0.27 -37% -49% 0.54
cpu_activity -45% 2% -89% 0.96 1.23 0.23 -20% -17% 0.78
cpu_activity -38% -1% -61% 0.96 1.29 0.24 -3% -16% 0.78
naval_propulsion_plant -36% -4% -52% 0.95 5.40 1.59 -215% -562% -0.85
naval_propulsion_plant -44% -2% -78% 0.95 4.14 1.22 -99% -94% 0.16
naval_propulsion_plant -64% -1% -175% 0.98 4.66 1.37 -141% -163% 0.50
naval_propulsion_plant -35% 2% -67% 0.92 3.53 1.04 -24% -27% 0.27
naval_propulsion_plant -38% -2% -70% 0.94 4.75 1.40 -62% -68% -0.21
miami_housing -32% -1% -46% 0.94 1.48 0.18 -21% -20% 0.74
miami_housing -46% 1% -88% 0.96 1.98 0.25 -41% -48% 0.53
miami_housing -25% -2% -31% 0.95 1.69 0.21 -18% -19% 0.75
miami_housing -34% -2% -49% 0.96 1.82 0.23 -23% -22% 0.77
miami_housing -42% -2% -72% 0.91 1.91 0.24 -24% -32% 0.57
kin8nm -14% 1% -17% 0.98 3.68 0.72 -3% 0% 0.43
kin8nm -18% 0% -22% 0.97 3.67 0.72 -20% -22% 0.36
kin8nm -17% 1% -22% 0.96 3.56 0.70 -12% -10% 0.39
kin8nm -24% 0% -32% 0.96 3.46 0.68 -26% -24% 0.36
kin8nm -20% 0% -25% 0.95 3.51 0.69 -20% -19% 0.31
concrete_compressive_strength -10% 2% -13% 0.95 2.80 0.57 -9% -6% 0.51
concrete_compressive_strength 9% -3% 8% 0.95 2.68 0.54 -15% -13% 0.59
concrete_compressive_strength -1% -4% 3% 0.95 2.86 0.58 -20% -23% 0.52
concrete_compressive_strength 16% -5% 16% 0.91 2.49 0.50 -10% -10% 0.46
concrete_compressive_strength 13% -3% 11% 0.96 2.40 0.48 -5% -9% 0.60
cars -23% -3% -33% 0.95 1.93 0.33 -3% -5% 0.75
cars -28% -2% -43% 0.97 1.94 0.33 -36% -38% 0.75
cars -27% -3% -39% 0.94 2.06 0.35 -34% -33% 0.70
cars -41% 6% -85% 0.97 2.70 0.46 -32% -37% 0.63
cars 8% -4% 10% 0.90 1.68 0.28 -32% -15% 0.48
energy_efficiency -12% 0% -14% 0.99 1.44 0.42 -8% 1% 0.97
energy_efficiency -4% -2% -2% 0.97 0.69 0.20 14% 18% 0.98
energy_efficiency -24% -3% -37% 0.94 1.23 0.36 8% 2% 0.92
energy_efficiency 10% -3% -1% 0.96 1.07 0.31 -8% 6% 0.96
energy_efficiency -7% -10% 0% 0.89 0.84 0.24 -1% -32% 0.93
california_housing -28% 1% -42% 0.95 3.00 0.71 -8% -9% 0.55
california_housing -20% -2% -25% 0.97 2.85 0.68 -1% 0% 0.68
california_housing -30% 0% -42% 0.96 2.90 0.69 -7% -3% 0.63
california_housing -16% -1% -19% 0.93 2.70 0.64 -4% -3% 0.54
california_housing -28% 0% -40% 0.96 2.83 0.67 -2% -1% 0.65
airfoil_self_noise -14% 0% -16% 1.00 3.46 0.74 -20% -11% 0.84
airfoil_self_noise 2% -6% 7% 0.89 2.60 0.56 -22% -13% 0.31
airfoil_self_noise -12% -8% -9% 0.88 2.56 0.55 -23% -20% 0.44
airfoil_self_noise -13% -5% -12% 0.93 2.94 0.63 -9% -1% 0.62
airfoil_self_noise -5% 1% -8% 0.94 2.77 0.59 -3% 1% 0.54
QSAR_fish_toxicity -16% 1% -20% 0.99 4.78 0.74 -8% -9% 0.44
QSAR_fish_toxicity -7% -5% -7% 0.92 3.31 0.51 -21% -28% 0.28
QSAR_fish_toxicity -5% -2% -8% 0.95 3.17 0.49 -6% -8% 0.39
QSAR_fish_toxicity -8% -5% -8% 0.86 2.74 0.43 -21% -17% 0.28
QSAR_fish_toxicity -17% 2% -28% 0.90 3.03 0.47 -10% -9% 0.32
Appendix F Aggregated results for non-conformalized predictions
Table 12: Test results for all models with complete-raw data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$. Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
MultiXGBs
space_ga -1%$\pm$56%$\pm$2-7%$\pm$60.89$\pm$0.01 0.20$\pm$0.01 -9%$\pm$1-10%$\pm$20.60$\pm$0.01
cpu_activity 25%$\pm$69%$\pm$112%$\pm$40.89$\pm$0.01 0.07$\pm$0.00 21%$\pm$1-1%$\pm$20.98$\pm$0.00
naval_propulsion_plant 156%$\pm$138%$\pm$158%$\pm$20.96$\pm$0.00 $0$14%$\pm$4-12%$\pm$20.99$\pm$0.00
miami_housing 61%$\pm$5-7%$\pm$243%$\pm$30.82$\pm$0.01 0.06$\pm$0.00 4%$\pm$43%$\pm$30.91$\pm$0.01
kin8nm 10%$\pm$35%$\pm$24%$\pm$40.90$\pm$0.01 0.38$\pm$0.01 -20%$\pm$1-22%$\pm$10.63$\pm$0.01
concrete_compressive_strength 13%$\pm$426%$\pm$6-11%$\pm$70.91$\pm$0.01 $0$26%$\pm$6-40%$\pm$110.85$\pm$0.01
cars 18%$\pm$815%$\pm$42%$\pm$80.89$\pm$0.02 0.12$\pm$0.01 -3%$\pm$3-1%$\pm$30.95$\pm$0.00
energy_efficiency 247%$\pm$8922%$\pm$660%$\pm$170.88$\pm$0.02 0.05$\pm$0.01 -15%$\pm$22-16%$\pm$171.00$\pm$0.00
california_housing 31%$\pm$50%$\pm$123%$\pm$40.88$\pm$0.01 $0$1%$\pm$1-1%$\pm$10.81$\pm$0.00
airfoil_self_noise 61%$\pm$414%$\pm$931%$\pm$160.81$\pm$0.07 0.19$\pm$0.04 -64%$\pm$32-73%$\pm$410.86$\pm$0.05
QSAR_fish_toxicity 10%$\pm$311%$\pm$3-1%$\pm$30.76$\pm$0.01 0.23$\pm$0.00 3%$\pm$31%$\pm$40.55$\pm$0.02
MultiETs
space_ga 9%$\pm$3-2%$\pm$19%$\pm$40.91$\pm$0.01 0.22$\pm$0.01 -16%$\pm$2-18%$\pm$30.54$\pm$0.02
cpu_activity 27%$\pm$3-8%$\pm$027%$\pm$20.90$\pm$0.00 $0$14%$\pm$1-20%$\pm$20.98$\pm$0.00
naval_propulsion_plant 219%$\pm$32-9%$\pm$071%$\pm$30.91$\pm$0.00 0.19$\pm$0.02 -320%$\pm$41-409%$\pm$360.96$\pm$0.01
miami_housing 69%$\pm$3-13%$\pm$048%$\pm$10.86$\pm$0.00 $0$10%$\pm$2-20%$\pm$10.90$\pm$0.00
kin8nm 10%$\pm$2-9%$\pm$017%$\pm$20.88$\pm$0.01 0.44$\pm$0.01 -35%$\pm$2-38%$\pm$20.48$\pm$0.01
concrete_compressive_strength 9%$\pm$9-12%$\pm$219%$\pm$80.81$\pm$0.01 0.24$\pm$0.02 -67%$\pm$7-94%$\pm$90.76$\pm$0.01
cars 11%$\pm$616%$\pm$2-5%$\pm$70.83$\pm$0.02 0.09$\pm$0.01 3%$\pm$31%$\pm$20.95$\pm$0.00
energy_efficiency -4%$\pm$50%$\pm$7-5%$\pm$140.65$\pm$0.03 0.02$\pm$0.00 3%$\pm$30%$\pm$11.00$\pm$0.00
california_housing 31%$\pm$2-7%$\pm$029%$\pm$10.90$\pm$0.00 0.41$\pm$0.01 -15%$\pm$1-21%$\pm$10.71$\pm$0.01
airfoil_self_noise 18%$\pm$25-15%$\pm$425%$\pm$170.83$\pm$0.04 0.26$\pm$0.06 -117%$\pm$45-141%$\pm$510.80$\pm$0.08
QSAR_fish_toxicity 12%$\pm$2-9%$\pm$119%$\pm$20.82$\pm$0.01 $0$6%$\pm$1-11%$\pm$10.53$\pm$0.01
MultiMLPs
space_ga 8%$\pm$10%$\pm$37%$\pm$30.81$\pm$0.01 0.14$\pm$0.00 0%$\pm$1-1%$\pm$10.75$\pm$0.01
cpu_activity 8%$\pm$3-9%$\pm$415%$\pm$60.72$\pm$0.02 0.05$\pm$0.00 7%$\pm$25%$\pm$20.98$\pm$0.00
naval_propulsion_plant 21%$\pm$180%$\pm$316%$\pm$130.92$\pm$0.01 0.07$\pm$0.01 -45%$\pm$25-36%$\pm$201.00$\pm$0.00
miami_housing 9%$\pm$42%$\pm$36%$\pm$60.81$\pm$0.02 $0$3%$\pm$24%$\pm$20.91$\pm$0.00
kin8nm 2%$\pm$78%$\pm$5-6%$\pm$100.81$\pm$0.02 0.13$\pm$0.01 7%$\pm$27%$\pm$20.93$\pm$0.00
concrete_compressive_strength 98%$\pm$13-23%$\pm$761%$\pm$50.56$\pm$0.04 0.06$\pm$0.00 12%$\pm$515%$\pm$40.91$\pm$0.01
cars 0%$\pm$146%$\pm$12-9%$\pm$300.76$\pm$0.08 0.07$\pm$0.01 -3%$\pm$30%$\pm$30.95$\pm$0.00
energy_efficiency 82%$\pm$1825%$\pm$2331%$\pm$110.62$\pm$0.04 0.02$\pm$0.00 32%$\pm$333%$\pm$31.00$\pm$0.00
california_housing -14%$\pm$28%$\pm$1-26%$\pm$30.89$\pm$0.01 0.31$\pm$0.01 7%$\pm$09%$\pm$10.82$\pm$0.00
airfoil_self_noise 74%$\pm$410%$\pm$639%$\pm$150.75$\pm$0.02 0.08$\pm$0.00 18%$\pm$416%$\pm$50.97$\pm$0.00
QSAR_fish_toxicity -6%$\pm$411%$\pm$3-19%$\pm$50.76$\pm$0.03 0.23$\pm$0.01 4%$\pm$57%$\pm$40.55$\pm$0.04
Table 13: Test results for all models with missing-raw data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$. Performance over the baseline is highlighted in bold.
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
MultiXGBs
space_ga -5%$\pm$64%$\pm$2-12%$\pm$70.86$\pm$0.02 0.22$\pm$0.01 -7%$\pm$4-9%$\pm$30.45$\pm$0.06
cpu_activity -5%$\pm$1912%$\pm$8-37%$\pm$460.89$\pm$0.06 0.14$\pm$0.04 -1%$\pm$21-17%$\pm$130.83$\pm$0.07
naval_propulsion_plant 105%$\pm$29-16%$\pm$656%$\pm$60.70$\pm$0.06 0.18$\pm$0.02 -14%$\pm$20-29%$\pm$350.77$\pm$0.05
miami_housing 16%$\pm$13-5%$\pm$510%$\pm$160.82$\pm$0.06 0.10$\pm$0.02 -14%$\pm$11-18%$\pm$100.72$\pm$0.10
kin8nm -5%$\pm$23%$\pm$2-14%$\pm$50.88$\pm$0.01 0.48$\pm$0.01 -3%$\pm$3-4%$\pm$40.40$\pm$0.02
concrete_compressive_strength 2%$\pm$38%$\pm$4-10%$\pm$30.77$\pm$0.03 0.30$\pm$0.01 -16%$\pm$12-20%$\pm$140.58$\pm$0.08
cars -15%$\pm$1426%$\pm$9-63%$\pm$260.85$\pm$0.04 0.27$\pm$0.03 -30%$\pm$22-35%$\pm$240.66$\pm$0.15
energy_efficiency 10%$\pm$2611%$\pm$6-24%$\pm$430.92$\pm$0.02 0.22$\pm$0.05 -186%$\pm$265-153%$\pm$1880.91$\pm$0.08
california_housing 26%$\pm$7-2%$\pm$321%$\pm$60.88$\pm$0.03 0.40$\pm$0.03 -4%$\pm$4-5%$\pm$50.69$\pm$0.05
airfoil_self_noise -2%$\pm$18-3%$\pm$4-4%$\pm$170.73$\pm$0.04 0.34$\pm$0.06 -37%$\pm$31-43%$\pm$360.41$\pm$0.12
QSAR_fish_toxicity -2%$\pm$30%$\pm$5-6%$\pm$70.73$\pm$0.03 0.26$\pm$0.02 -4%$\pm$4-2%$\pm$30.36$\pm$0.04
MultiETs
space_ga 4%$\pm$3-5%$\pm$17%$\pm$40.88$\pm$0.02 0.24$\pm$0.01 -12%$\pm$4-14%$\pm$40.40$\pm$0.05
cpu_activity -16%$\pm$10-5%$\pm$1-20%$\pm$180.93$\pm$0.01 0.20$\pm$0.03 -30%$\pm$23-54%$\pm$200.79$\pm$0.10
naval_propulsion_plant 58%$\pm$5-17%$\pm$446%$\pm$30.82$\pm$0.05 0.41$\pm$0.03 -36%$\pm$21-93%$\pm$480.71$\pm$0.05
miami_housing 20%$\pm$14-10%$\pm$421%$\pm$130.87$\pm$0.05 0.14$\pm$0.03 -27%$\pm$11-37%$\pm$110.67$\pm$0.11
kin8nm 5%$\pm$2-8%$\pm$212%$\pm$40.87$\pm$0.02 0.50$\pm$0.02 -11%$\pm$2-13%$\pm$30.32$\pm$0.03
concrete_compressive_strength 1%$\pm$5-14%$\pm$313%$\pm$20.80$\pm$0.02 0.37$\pm$0.01 -35%$\pm$21-50%$\pm$250.49$\pm$0.06
cars -23%$\pm$6-8%$\pm$5-35%$\pm$140.83$\pm$0.04 0.24$\pm$0.02 -33%$\pm$18-31%$\pm$180.66$\pm$0.09
energy_efficiency -15%$\pm$160%$\pm$9-36%$\pm$420.85$\pm$0.05 0.15$\pm$0.01 -73%$\pm$74-84%$\pm$640.94$\pm$0.04
california_housing 21%$\pm$7-7%$\pm$222%$\pm$50.91$\pm$0.02 0.52$\pm$0.03 -14%$\pm$3-17%$\pm$30.61$\pm$0.05
airfoil_self_noise -7%$\pm$18-20%$\pm$28%$\pm$160.76$\pm$0.03 0.42$\pm$0.08 -65%$\pm$54-93%$\pm$700.33$\pm$0.11
QSAR_fish_toxicity 10%$\pm$7-9%$\pm$316%$\pm$30.82$\pm$0.03 0.33$\pm$0.03 -8%$\pm$8-10%$\pm$90.39$\pm$0.09
MultiMLPs
space_ga -11%$\pm$5-15%$\pm$21%$\pm$60.68$\pm$0.03 0.16$\pm$0.01 -40%$\pm$23-34%$\pm$150.21$\pm$0.23
cpu_activity -50%$\pm$1622%$\pm$14-193%$\pm$1180.86$\pm$0.10 0.18$\pm$0.07 -15%$\pm$19-23%$\pm$170.75$\pm$0.12
naval_propulsion_plant -52%$\pm$19-14%$\pm$24-102%$\pm$560.63$\pm$0.18 0.34$\pm$0.07 -108%$\pm$66-183%$\pm$195-0.03$\pm$0.47
miami_housing -33%$\pm$1513%$\pm$10-92%$\pm$700.85$\pm$0.06 0.12$\pm$0.04 -25%$\pm$8-28%$\pm$110.67$\pm$0.10
kin8nm -11%$\pm$5-5%$\pm$2-8%$\pm$60.77$\pm$0.01 0.35$\pm$0.02 -16%$\pm$8-15%$\pm$90.37$\pm$0.04
concrete_compressive_strength -15%$\pm$13-13%$\pm$28-10%$\pm$480.60$\pm$0.22 0.22$\pm$0.11 -12%$\pm$5-12%$\pm$60.54$\pm$0.05
cars -51%$\pm$443%$\pm$7-211%$\pm$340.93$\pm$0.02 0.29$\pm$0.03 -28%$\pm$12-26%$\pm$130.66$\pm$0.10
energy_efficiency -13%$\pm$189%$\pm$15-30%$\pm$250.66$\pm$0.13 0.09$\pm$0.04 1%$\pm$9-1%$\pm$160.95$\pm$0.03
california_housing -17%$\pm$64%$\pm$3-28%$\pm$140.88$\pm$0.03 0.44$\pm$0.04 -5%$\pm$2-3%$\pm$30.61$\pm$0.06
airfoil_self_noise 2%$\pm$16-2%$\pm$5-1%$\pm$210.75$\pm$0.07 0.25$\pm$0.08 -15%$\pm$8-9%$\pm$80.55$\pm$0.18
QSAR_fish_toxicity -17%$\pm$56%$\pm$6-30%$\pm$110.76$\pm$0.04 0.28$\pm$0.03 -13%$\pm$7-14%$\pm$80.34$\pm$0.06
Appendix G Example of learning with non-normality